Two charges lie on positive x-axis. Charge A (5.0 x 10-9 C) is 4.0 cm from the origin and Charge B is 6.0 cm from the origin. (-7.0 x 10-9 C). What is the total force exerted by these two charges on Charge C (8.0 x 10-9 C) located at the origin?


1
Expert's answer
2022-03-08T08:56:43-0500

The net force

"F=F_{AC}-F_{BC}=k\frac{q_Aq_C}{r_1^2}-k\frac{q_Bq_C}{r_2^2}""F=9*10^9\left(\frac{5*10^{-9}*8*10^{-9}}{0.04^2}-\frac{7*10^{-9}*8*10^{-9}}{0.06^2}\right)\\
=8.5*10^{-5}\:\rm N"